### v3

# Interaction-induced reentrance of Bose glass and quench dynamics of Bose gases in twisted bilayer and quasicrystal optical lattices

Abstract
We investigate the ground state and dynamical properties of ultracold gases in optical lattices with a quasicrystal structure—a scenario inspired by recent experiments on twisted bilayer optical lattices and optical quasicrystals. Our study reveals that the interplay between on-site repulsive interactions and a quasiperiodic potential gives rise to rich physics. At low filling factors, increasing the interaction strength induces a delocalization effect that transforms a Bose-glass (BG) phase, characterized by disconnected superfluid (SF) regions, into a robust SF phase with a percolated network of SF clusters. This transition is quantitatively characterized by monitoring the percolation probability. At higher filling factors, we uncover a striking reentrant behavior: As the on-site interaction increases, the system initially transitions from BG to SF, but a further increase reverses this trend, returning the system to the BG phase. This reentrance is ascribed to an interaction-driven rearrangement of particles, where a once percolated SF network fragments into isolated SF islands as repulsive interactions dominate. Furthermore, our analysis of quench dynamics demonstrates distinct transient behaviors. Intra-phase quenches yield minimal variations in both the percolation probability and the inverse participation ratio (IPR) of the particle density distribution. In contrast, interphase quenches produce pronounced effects; for instance, a quench from the SF to BG phase is marked by an abrupt loss of global SF connectivity, while a BG-to-SF quench features oscillatory changes in the percolation probability and a gradual decrease in the IPR, eventually stabilizing the SF phase. Our findings unveil the complex interplay between quasiperiodic optical lattice potential and interaction in ultracold Bose gases, offering valuable insights that are highly pertinent to current experimental efforts employing state-of-the-art twisted bilayer and quasicrystalline optical lattice platforms.
I Introduction
Recently, twisted bilayer materials have opened a brand-new era of twistronics, which has become fertile ground for studying and engineering strong-correlation phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Most research has focused on moiré patterns, the commensurate patterns of multiple twisted crystal layers, in which the whole system retains certain translational symmetry but the unit cell is much larger than the original one, leading to the flattening of bands and thus enhancing interaction effects. Experimentalists have observed in the laboratory many emergent strong-correlation effects, such as correlated insulators [1, 2], superconductivity [3], magnetism [4, 5, 6], generalized Wigner crystals [7, 8], and integer and fractional quantum anomalous Hall insulators [9, 10, 11, 12, 13, 14, 15, 16].
On the other hand, quasicrystal patterns can emerge when the crystal layers are twisted in an incommensurate way that completely breaks the translational symmetry but preserves the long-range order [17], thereby mediating between order and disorder. Quasicrystals have been experimentally realized and highly tuned by twisting three layers of graphene [18] or bilayers of tungsten diselenides [19] but have received less attention before [20, 21] than the commensurate case. It is well-known that a quasicrystal will lead to Anderson localization [22], a phenomenon that lacks the diffusiveness of noninteracting particles when it is subjected to a random or quasiperiodic potential landscape. In the one-dimensional (1D) and two-dimensional (2D) cases, all single-particle eigenstates can be localized by a finite strength of the quasicrystal potential [23, 24, 25], different from the random disorder where the localization occurs for any infinitesimal strength [26]. Therefore, quasicrystals are ideal platforms for studying localization physics.
Now that the moiré pattern can induce many exotic strongly correlated phenomena, one may wonder how the localization feature in twisted bilayer quasicrystals competes with the interaction. In fermionic systems, most research has focused on the thermal properties of many-body localization, especially in the 1D case [27, 28, 29, 30, 31]. The incommensurability is also found to radically change the ground state of interacting fermions, compared with the commensurate case [32, 33, 34, 35, 36, 37]. In bosonic systems, the interplay of the randomness or quasiperiodicity and the interaction will lead to an exotic many-body ground state—Bose glass (BG) [38, 39], which is insulating like an Anderson insulator but compressible like a superfluid. Disordered interacting bosons in different dimensions have been studied using various experimental platforms, such as helium films on porous substrates [40], disordered superconducting films [41], doped quantum magnets [42], and, especially, optical lattices [43, 44, 45, 46, 47, 48]. There are also plenty of theoretical studies on interacting bosons in disordered or quasiperiodic potentials [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64] and even bosons in interaction-induced disordered or quasiperiodic potentials [65]. The experimental realization of twisted bilayer optical lattices [66] has opened new avenues to investigate interacting bosons in quasicrystals using concrete experimental platforms [67, 68]. More recently, 2D octagonal quasicrystalline optical lattices were directly realized in ultracold-atom experiments by using four light beams in a plane and one beam in the perpendicular direction, and BG with ultracold atoms in the weakly interacting regime was observed [69].
In this work, we study the ground state and dynamical properties of a class of ultracold bosonic systems realized in optical lattices with an aperiodic external potential—a setting directly inspired by recent experiments on twisted bilayer optical lattices [66] and optical quasicrystals [69]. We investigate how the interplay between on-site repulsive interactions and a quasiperiodic potential drives the system between a homogeneous superfluid (SF) phase and a BG phase. At low filling factors, our findings reveal that an increase in the on-site interaction strength induces a delocalization effect, transforming a BG phase—with disconnected SF regions lacking global connectivity—into a robust SF phase characterized by a percolated network of SF clusters. This phase transition is quantitatively characterized by monitoring the percolation probability (see Fig. 1). At higher filling factors, our results uncover a striking reentrant behavior (see Fig. 2): The system initially undergoes a transition from a BG to an SF phase as the on-site interaction strength increases, but a further increase in the interaction leads to a reversal back to the BG phase. This reentrance is attributed to the interaction-driven rearrangement of particles, wherein a once-percolated SF network eventually fragments into isolated SF islands as repulsive interactions become dominant. Moreover, our examination of the system’s quench dynamics reveals distinct transient behaviors (see Fig. 3): While intraphase quenches produce minimal changes in both the percolation probability and inverse participation ratio (IPR) of the particle-density distribution, interphase quenches lead to pronounced dynamical effects. Specifically, a quench from the SF phase to the BG phase results in an abrupt loss of global SF connectivity, whereas a BG-to-SF quench is marked by oscillatory changes in the percolation probability and a gradual decrease in the IPR, ultimately stabilizing the SF phase. These findings unveil the rich physics of ultracold Bose gases in optical lattices with an aperiodic external potential, offering valuable insights that are highly relevant for state-of-the-art experiments employing twisted bilayer or quasicrystalline optical lattice platforms.
II System and Model
Inspired by the recent experimental progress in realizing twisted bilayer optical lattice [66] and optical quasicrystal [69] for ultracold atoms, we consider a single-component ultracold bosonic system, the Hamiltonian of which takes the form of a “single-layer” Bose-Hubbard model with an additional external on-site potential, i.e.,
$\hat{H}=-J\sum_{\left\langle i,j\right\rangle}\hat{b}_{i}^{\dagger}\hat{b}_{j}+\frac{U}{2}\sum_{i}\hat{n}_{i}\left(\hat{n}_{i}-1\right)+\sum_{i}M_{i}\hat{n}_{i},$ (1)
where $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$ is the occupation number operator at site $i$, with $\hat{b}_{i}^{\dagger}\penalty 10000\ (\hat{b}_{i})$ being the corresponding bosonic creation (annihilation) operator; $U$ is the strength of the on-site interaction; and $J$ is the hopping amplitude between nearest-neighbor lattice sites (denoted by $\langle\cdots\rangle$). Here, we focus on the case where the underlying optical lattice is a square lattice because it is the most relevant for the current experimental setups [66, 69]. The last term of $\hat{H}$ is the additional on-site external potential with $M_{i}$ being its strength at site $i$. In the experimental setup that realizes the twisted bilayer optical lattice [66], the on-site external potential $M_{i}$ takes into account the influence of the lattice on the other layer in the large-detuning limit, where the bilayer system can be effectively described by the single-layer model [70, 66]. While in the experimental set-up that realizes the optical quasicrystal [69], $M_{i}$ simply corresponds to the additional optical lattice potential felt by the atoms. In this regard, the physics of the model Hamiltonian (1) that will be discussed in the following is relevant for ultracold twisted bilayer optical lattice systems [66] and optical quasicrystal systems [69] in experiments.
In the following, we focus on the case where $M_{i}$ assumes the form
$M_{i}=$ $M_{r}\left[\sin^{2}\left(i_{x}\pi\cos\theta+i_{y}\pi\sin\theta\right)\right$.
$\left.+\sin^{2}\left(i_{y}\pi\cos\theta-i_{x}\pi\sin\theta\right)\right],$ (2)
where $M_{r}$ is the strength of the external potential, $i_{x}$ ($i_{y}$) is the index of the $i\textrm{th}$ site of the square lattice along the $x$ ($y$) direction, and we choose $\theta=45\text{\textdegree}$. This corresponds to the case in bilayer systems with a twist angle of $45\text{\textdegree}$ [66] or the one in the optical quasicrystal system with the optical lattice formed by superimposing four independent one-dimensional lattices in the $x$-$y$ plane at a $45\text{\textdegree}$ angle [69].
The first two terms of the Hamiltonian (1) assume the form of the conventional Bose-Hubbard model [39]; they favor two homogeneous phases, namely, the homogeneous SF phase (for large $J/U$) and the Mott insulator phase (for small $J/U$ at integer filling factors) that respect the discrete translation symmetry of the underlying lattice. The third term of $\hat{H}$ is an aperiodic or a quasiperiodic external potential that explicitly breaks the lattice translation symmetry. Recent investigations [60] showed that this type of term can give rise to the BG phase that usually arises in disordered bosonic systems [39] due to their aperiodicity. In fact, the interplay between on-site interaction and the quasiperiodic external potential can give rise to much richer physics associated with the BG phase. Particularly, at fixed particle densities that are highly relevant for corresponding experiments, this interplay can give rise to interaction induced reentrant transitions for BG phases and rich quench dynamics as we shall show in the following.
III Results
To study the ground state of the system described by Hamiltonian (1), we used the bosonic Gutzwiller variational method [71, 72, 73]. The variational ground state is assumed to take the site-factorized form $\left|\textrm{GW}\right\rangle=\left|\phi_{1}\right\rangle\otimes\cdots\otimes\left|\phi_{i}\right\rangle\otimes\cdots\otimes\left|\phi_{N_{\textrm{lat}}}\right\rangle$, where each $\left|\phi_{i}\right\rangle$ represents the state at the $i$th lattice site and $N_{\textrm{lat}}$ denotes the total number of sites in the lattice. The individual single-site states $\left|\phi_{i}\right\rangle$ are expanded as a superposition of particle-number eigenstates as $\left|\phi_{i}\right\rangle=\sum_{n=0}^{n_{\textrm{max}}}C_{i.n}\left|n\right\rangle_{i}$, where $n_{\textrm{max}}$ is the maximum number of particles per site and $C_{i.n}$ are variational parameters that are to be determined by minimizing the total energy of the system. In this work, we focus on the ground-state properties of the system at different filling factors $\rho\equiv N/N_{\textrm{lat}}$, with $N$ being the total number of particles in the system. We examine the system at different filling factors to explore how particle density influences the ground-state properties. For the numerical calculations, we set square-lattice size $L=51$, the local maximum truncation number $n_{\textrm{max}}=10$, and we employ open boundary conditions for the numerical results if not otherwise specified in the text.
Figure 1: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $100$ particles on a $51\times 51$ square lattice, where the black solid line is the phase boundary between BG (top) and SF (bottom) phases. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and the inverse participation ratio (IPR; red curve) on $U/J$ for points along the dashed line indicated in the phase diagram. The inserted SF order-parameter distributions correspond to the nearby red points marked in the main plot. (b) Real-space distributions of the particle density $\left\langle\hat{n}_{i}\right\rangle$, the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, and the discrete field $S_{i}$ near the phase boundary ($M_{r}/J=5.9$). See text for more details.
III.1 Transitions between superfluids and Bose glasses
For relatively low filling factors, the typical properties of the system are summarized in Fig. 1, which shows the phase diagram of the system at $\rho\approx 0.04$, along with the particle-density distribution $\left\langle\hat{n}_{i}\right\rangle$, the SF order parameter distribution $\langle\hat{b}_{i}\rangle$, and the discrete field distribution $S_{i}$. The latter is used to distinguish between SF and BG phases and will be discussed in detail later. As shown in the bottom left inset of Fig. 1(a), in the low-aperiodic-potential regime, as expected from the homogeneous Bose-Hubbard model, the system is in the SF phase, with a nonzero SF order parameter $\langle\hat{b}_{i}\rangle$ that is distributed relatively evenly across the entire lattice. In contrast, in the high-aperiodic-potential regime, as shown in the top inset of Fig. 1(a), the system maintains a nonzero SF order parameter $\langle\hat{b}_{i}\rangle$, but its distribution exhibits a qualitatively different structure. In particular, although the SF order parameter is nonzero in various regions, these regions are “disconnected” from each other, indicating a transition to a new phase: the BG phase, which is absent in conventional homogeneous systems of lattice bosons.
To quantitatively distinguish the SF phase from the BG phase, we employ the percolation method [74, 75, 76, 77, 60], a widely used technique for studying phase transitions in disordered or quasidisordered systems. The key quantity used to differentiate the SF and BG phases is the percolation probability $\mathcal{P}$, with $\mathcal{P}=0$ indicating no connected SF region spanning the entire system (signifying the BG phase) and $\mathcal{P}>0$ indicating the presence of a percolated SF region (signifying a SF phase spanning the entire system). In this work, we adopt the method introduced in [77, 60], where $\mathcal{P}$ is calculated using the corresponding discrete field $S_{i}$ [77], defined as
$S_{i}=\begin{cases}0&\textrm{if}\textrm{ }I-\gamma_{n}\leq\left\langle\hat{n}_{i}\right\rangle\leq I+\gamma_{n},\\ 1&\textrm{otherwise},\end{cases}$ (3)
where $\gamma_{n}=5\times 10^{-3}$ is the threshold used in the numerical calculations and $I=0,1,2,\ldots$ denotes non-negative integer values. This discrete field encapsulates the percolation properties of the SF regions in the system (see Appendix A for technical details on extracting $\mathcal{P}$ from $S_{i}$). For example, in the last row of Fig. 1(b), red sites represent various connected SF regions that do not span the entire system, while blue sites correspond to connected SF regions that do span the entire system, forming a percolated SF region.
From Fig. 1(a), we observe that tuning the on-site interaction strength $U$ can drive the transition between the SF and BG phases in the high-aperiodic-potential regime. As indicated by the horizontal arrow in Fig. 1(a) and the corresponding distributions for $\left\langle\hat{n}_{i}\right\rangle$, $\langle\hat{b}_{i}\rangle$, and $S_{i}$ in Fig. 1(b) at a fixed aperiodic potential strength $M_{r}$, the system transitions from the BG phase with $\mathcal{P}=0$ at relatively small $U$ to the SF phase with a finite percolation probability $\mathcal{P}$ at larger $U$ [see the bottom right inset of Fig. 1(a)]. This transition is driven by a “delocalization” effect of the repulsive on-site interaction: Under low-filling conditions ($\rho<1$), bosons redistribute from highly occupied sites to nearby vacant or weakly occupied sites, broadening the density profile. By contrast, at integer filling, the standard localization effect of strong interactions can also manifest in this system, driving it into a Mott insulator (see Fig. 5 and Appendix B for details). As seen in the bottom right inset of Fig. 1(a), the IPR for the normalized density distribution $\widetilde{\left\langle\hat{n}_{i}\right\rangle}\equiv\left\langle\hat{n}_{i}\right\rangle/\sum_{i=1}^{N_{\textrm{lat}}}\left\langle\hat{n}_{i}\right\rangle$ directly characterizes the extent of localization in the system, i.e., $\textrm{IPR}\equiv\sum_{i=1}^{N_{\textrm{lat}}}\widetilde{\left\langle\hat{n}_{i}\right\rangle}^{2}$. The IPR decreases as $U$ increases, reflecting the delocalization of the particles. At relatively small $U$, the SF regions are localized [see the left most plot for $S_{i}$ in Fig. 1(b)]. As $U$ increases, the SF regions expand, and at large enough $U$, they can merge, forming a percolated SF region that drives the system into the SF phase.
Figure 2: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $196$ particles on a $51\times 51$ square lattice, where the black solid line is the phase boundary between BG (upper) and SF (lower) phases. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and IPR (red curve) on $U/J$ for points along the dashed line indicated in the phase diagram. (b) Real-space distributions of the particle density $\left\langle\hat{n}_{i}\right\rangle$, the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, and the discrete field $S_{i}$ near the phase boundary ($M_{r}/J=8$). See text for more details.
III.2 Reentrance between superfluid and Bose glass phases
In the previous section, we discussed the system at relatively low filling factors. From the particle density distribution in Fig. 1(b), we observe that the average number of particles per lattice site is less than 1, indicating the influence of the on-site interaction is very weak. This thus raises the question of whether different physical phenomena might arise when the system’s filling factor is higher. To explore this, we present the results for a system with a filling factor $\rho=0.08$, and the corresponding results are summarized in Fig. 2, which includes the phase diagram and typical real-space distributions. Comparing the results to the phase diagram at low filling factors, we observe a reentrant transition at higher filling factors, as shown in Fig. 2(a). Focusing on the dashed line in this phase diagram, we see that as the on-site interaction $U$ increases, the system first transitions from the BG phase to the SF phase. However, as $U$ continues to increase, the system transitions back from the SF phase to the BG phase. This indicates that the on-site interaction can drive a reentrant transition to the BG phase.
To understand this reentrant behavior, let us first check the typical real-space distributions of the system as $U$ changes shown in Fig. 2(b). When the on-site interaction $U/J=1.25$ is relatively small compared to the external potential field strength $M_{r}/J=8$, the system consists of isolated SF clusters. As we gradually increase the on-site interaction to $U/J=1.75$, a percolated SF region (marked in blue) appears in the $S_{i}$ distribution shown in Fig. 2(b), indicating that the system has transitioned from the BG phase to the SF phase.
The mechanism of this stage of transition is actually similar to that of the lower-filling case presented in the previous section. Specifically, when the on-site interaction $U$ increases, the most significant interaction effect occurs at lattice sites where the average number of particles is greater than 1. At these sites, the particles repel each other, leading to a redistribution. As a result, SF paths emerge during this repulsion process, and the system transitions from the BG phase to the SF phase.
However, as we further increase $U/J=3$, the system exhibits behavior different from that at low filling factors. As shown in the bottom right plot of Fig. 2(b), the percolated SF region disappears from the distribution of $S_{i}$ at $U/J=3$, indicating the system transitions from the SF phase back to the BG phase again. To understand this stage of transition, we note that although percolated SF regions can form in the SF phase, the increasing repulsive interaction strength can still cause the particle density distribution to undergo further rearrangement, causing percolated SF regions to break up into isolated SF islands, thus driving the system to enter the BG phase again. Importantly, this transition is characterized by the loss of global connectivity of the SF regions, while the local-density distribution changes only slightly. As a result, the IPR remains small [see the inset of Fig. 2(a)], even though the percolation probability $\mathcal{P}$ drops to zero, signaling the loss of long-range coherence and the onset of the BG phase. Moreover, if the particle density of each site is sufficiently low, that would indicate the local state of each site consists of only a zero- or single-occupation state. As a consequence, further increasing on-site interaction strength can hardly impose any physical influence on the system, consistent with the behavior seen in the last three columns of Fig. 1(b) for the low-filling case. In fact, at even high filling factors, one can observe even more times of reentrance (see Appendix B for more numerical results at even higher filling factors).
Figure 3: (a) Phase diagram of the system at $N=100$ with quench directions marked with arrows. (b) Time dependence of the IPR of the corresponding quench dynamics. (c) Time dependence of the percolation probability $\mathcal{P}$ of the corresponding quench dynamics. Representative real-space distributions of the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, the phase $\theta_{i}\equiv\arg\langle\hat{b}_{i}\rangle$ of the SF order parameter, local winding number $w_{i}$ of the phase field $\theta_{i}$, and the discrete field $S_{i}$ during (d) intraphase and (e) interphase quench processes indicated by the arrows in (a). See text for more details.
III.3 Quench dynamics
Thus far, we have focused on the static properties of the system. Now let us turn to the dynamical properties; in particular, here we shall focus on the quench dynamics of the system. Specifically, via the time-dependent Gutzwiller approach [78, 79, 80] (see Appendix C for technique details of the approach), we investigate quench dynamics of two different types, namely, intraphase quench and interphase quench.
Figure 3(d) shows the representative distributions of the system during the intraphase quenches, where the system remains within the same phase (e.g., from the BG or SF phase to the BG or SF phase). Neither the time evolution of the IPR nor the percolation probability $\mathcal{P}$ exhibits significant changes [see Figs. 3(b) and 3(c)]. Likewise, the typical real-space distributions of the amplitude of the SF order parameter and the discrete field at different time instants show no significant variations [see Fig. 3(d)]. However, we do notice the distribution of the local winding number $w_{i}$, defined for each plaquette of the lattice as the winding of the phase field $\theta_{i}\equiv\arg\langle\hat{b}_{i}\rangle$ associated with the SF order parameter, i.e., $w_{i}\equiv[(\theta_{i_{x}+1,i_{y}}-\theta_{i_{x},i_{y}})+(\theta_{i_{x}+1,i_{y}+1}-\theta_{i_{x}+1,i_{y}})+(\theta_{i_{x},i_{y}+1}-\theta_{i_{x}+1,i_{y}+1})+(\theta_{i_{x},i_{y}}-\theta_{i_{x},i_{y}+1})]/2\pi$, undergoes noticeable changes during the BG to BG quench. In this case, vortex pairs with opposite charges, indicated by the nonzero local winding number $w_{i}$, exhibit creation and annihilation dynamics. This vortex behavior can be attributed to the reorganization of the particle density distribution of the system during the quench dynamics, in which the locations of sites with vanishing particle density—fertile ground for the creation of vortex pairs—vary over time.
For interphase quenches, in which the system undergoes a transition between different phases (e.g., from SF to BG or vice versa), we observe more significant qualitative changes. Specifically, for these quenches, the initial state corresponds to the ground-state configuration at the two red dots in Fig. 3(a) with $(U/J=2,M_{r}/J=5)$ in the SF phase and $(U/J=2,M_{r}/J=10)$ in the BG phase. In the quench from the SF phase to the BG phase, the percolation probability $\mathcal{P}$ [illustrated by the green line in Fig. 3(c)] rapidly decreases from nearly 1 to 0, indicating the swift transition of the system from SF to BG. This sharp drop in $\mathcal{P}$ originates from the fragmentation of the dominant percolated SF cluster into disconnected SF islands, a process driven by particle hopping in the vicinity of the percolated path; accordingly, the timescale of this drop is determined by the hopping amplitude. Once the system reaches the BG phase, $\mathcal{P}$ remains at zero, reflecting that the system remains in the BG phase. This behavior arises because the BG phase consists of isolated SF islands, which restrict particle movement to these disconnected regions. As a result, the IPR remains relatively constant with only minor fluctuations, corresponding to slight variations in the spatial distribution of particles over time [see the green line in Fig. 3(b)]. It is worth noting that the IPR does not fully rise to the value corresponding to the final BG ground state. This difference originates from the fact that the late-time state after the SF-to-BG quench is a nonequilibrium state evolving from the SF ground state, whose conserved energy differs from that of the BG ground state at the same parameters. The local-winding-number distribution $w_{i}$ also exhibits rapid dynamics following the quench, characterized by an explosive generation of vortex pairs [see the right panel in Fig. 3(e)]. This behavior signifies the breakdown of phase coherence on progressively smaller length scales.
For the quench from the BG phase to the SF phase, the percolation probability $\mathcal{P}$ fluctuates between zero and nonzero values for times $tJ>5$, signaling the system’s transition between the BG and SF phases [see the blue line in Fig. 3(c)]. Unlike the quench from SF to BG, in which particles are confined to isolated SF islands, in the BG-to-SF quench, particles can move freely between SF clusters, allowing the formation of SF paths. However, at certain moments, the system may lose its SF path and transition back to the BG phase. This intermittent behavior gradually fades over time, and as shown in Fig. 3, the periods during which $\mathcal{P}=0$ become shorter, eventually leading the system to settle into a steady SF phase. This trend is also reflected in the time evolution of the IPR in Fig. 3(b). Initially, in the BG phase, the particles are localized at specific lattice sites, resulting in a relatively large IPR. As the system evolves, particles diffuse through space, causing a gradual decrease in the IPR. Although the IPR may occasionally increase, it generally decreases over time, indicating that the system eventually transitions to the SF phase. A similar pattern appears in the dynamics of the local winding number distribution $w_{i}$ [see the left panel in Fig. 3(e)]. Over time, the number of vortex pairs generally decreases, indicating the system’s progressive restoration of phase coherence across increasingly larger length scales.
IV CONCLUSIONS
Our work demonstrated that ultracold bosonic systems in optical lattices with an aperiodic external potential exhibit a rich interplay between disorder and interaction effects. We showed that at low filling factors, an increase in the on-site repulsive interaction triggers a delocalization process that transforms a BG phase, characterized by isolated and disconnected SF regions, into a robust SF phase with a percolated network. This transition is clearly evidenced by the evolution of the percolation probability and IPR. At higher filling factors, we uncovered a striking reentrant behavior wherein the system initially transitions from a BG to a SF phase as the interaction strength increases, but further enhancement of the interaction causes the SF network to fragment, causing the system to revert back to a BG phase. Moreover, our analysis of quench dynamics revealed markedly different transient behaviors for intraphase and interphase quenches, highlighting the sensitive dependence of the system’s dynamical evolution on the initial state and interaction parameters. These findings not only deepen our understanding of localization phenomena in interacting Bose gases but also offer valuable insights for ongoing experiments utilizing twisted bilayer and quasicrystalline optical lattice platforms, where tuning disorder and interaction effects is pivotal for exploring novel many-body phases.
Acknowledgements.
This work is supported by the NKRDPC (Grant No. 2022YFA1405304), the NSFC (Grant No. 12275089, No. 12574193, and No. 11904109), the Guangdong Basic and Applied Basic Research Foundation (Grants No. 2024A1515010188, No. 2023A1515012800, and No. 2021A1515010212), the Guangdong Provincial Key Laboratory (Grant No. 2020B1212060066), and the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2401002).
V Data Availability
The data that support the findings of this article are openly available [81].
Appendix A TECHNICAL DETAILS FOR CALCULATING PERCOLATION PROBABILITY FROM A DISCRETE FIELD
In this appendix, we provide technical details for calculating the percolation probability $\mathcal{P}$ from the discrete field $S_{i}$. The percolation probability $\mathcal{P}$ is defined in Ref. [60] as
$\mathcal{P}=\frac{N_{\textrm{span}}}{N_{\varphi}},$ (4)
where $N_{\textrm{span}}$ is the number of SF sites in a percolating cluster and $N_{\varphi}$ is the total number of sites with finite $\varphi$. In our numerical calculations, $N_{\textrm{span}}$ and $N_{\varphi}$ are derived from the discrete field $S_{i}$ [77], which can be constructed by the particle density distribution $\left\langle\hat{n}_{i}\right\rangle$ and SF order-parameter distribution $\langle\hat{b}_{i}\rangle$ described in the main text. According to the definition of $S_{i}$ [see Eq. (3)], $N_{\varphi}$ corresponds to the total number of sites where $S_{i}=1$, i.e., $N_{\varphi}=\sum_{i}^{N_{\textrm{lat}}}S_{i}$. To calculate $N_{\textrm{span}}$, it is necessary to identify the percolating cluster. The criterion for percolation, as outlined in [82], is the existence of at least one SF path that spans the entire system, either from left to right or from top to bottom. Once such an SF path is identified, the SF cluster that contains this SF path is considered the percolating cluster. The total number of lattice sites in this percolating cluster is then defined as $N_{\textrm{span}}$. Clearly, as shown by the calculation of the percolation probability $\mathcal{P}$ described above, this quantity can be used to distinguish the SF phase from the BG phase. In the BG phase, there are no percolating SF clusters, which results in a percolation probability of zero. In contrast, in the SF phase, the presence of percolating SF clusters leads to a nonzero percolation probability.
Appendix B NUMERICAL RESULTS AT EVEN HIGHER FILLING FACTORS
As discussed in the main text, we hypothesize that multiple reentrance phenomena could be observed in the system with an even higher particle filling. Indeed, as anticipated, this behavior is demonstrated in Fig. 4.
Figure 4: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $1024$ particles on a $51\times 51$ square lattice. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and IPR (red curve) on $U/J$ for points along the dashed line indicated in the phase diagram. (b) Real-space distributions of the particle density distribution $\left\langle\hat{n}_{i}\right\rangle$, the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, and the discrete field $S_{i}$ of four points along the dashed line ($M_{r}/J=25$). See text for more details.
We plot the phase diagram for a particle filling of 1024, which reveals two distinct peaks. Additionally, we includ a series of real-space distribution diagrams showing the effect of varying the interaction strength $U$ on the lattice, with the changes in $U$ indicated by the horizontal lines in the phase diagram. As shown in the inset of the phase diagram, the percolation probability of the ground state exhibits multiple transitions between zero and nonzero values, corresponding to the system’s behavior as it transitions from the BG phase to the SF phase, then back to the BG phase, and finally returns to the SF phase. The particle-density-distribution diagram Fig. 4(b) further illustrates that, as the particle filling increases, the system undergoes more significant changes due to the enhanced on-site interaction strength. Moreover, these changes lead to the formation and annihilation of SF path, creating a more complex phase transition process.
Figure 5: Dependence of the spatially averaged SF order parameter $\sum_{i}|\langle\hat{b}_{i}\rangle|/N_{\mathrm{lat}}$ on the interaction strength at different filling factors. The aperiodic potential strength is fixed at $M_{r}=7$, and the lattice size is $L=51$ ($N_{\mathrm{lat}}=51\times 51=2601$).
Moreover, at integer fillings, the standard localization effect of strong interactions can also manifest in this system, driving it into a Mott insulator phase. Figure 5 shows the dependence of the spatially averaged superfluid (SF) order parameter $\sum_{i}|\langle\hat{b}_{i}\rangle|/N_{\mathrm{lat}}$, on the interaction strength $U$ for the different fillings. At integer filling $\rho=1$, the SF order parameter vanishes at sufficiently large $U/J$, clearly signaling the emergence of a Mott insulator phase.
Appendix C Time dependent Gutzwiller approach
In this appendix, we present the technical details of the time-dependent Gutzwiller approach [78, 79, 80]. This approach is built on top of the static Gutzwiller approach. The static Gutzwiller method is employed to determine the ground state by minimizing the total energy of the system, given by $E\left(\left\{C_{i,n}\right\}\right)=\left\langle\textrm{GW}\right|\hat{H}\left|\textrm{GW}\right\rangle$, while the dynamic Gutzwiller method minimizes the action to simulate the system’s evolution. Specifically, by applying the variational principle, we can write the action of the system as $S=\int dt\left[-i\left\langle\textrm{GW}\left(t\right)\right|\partial_{t}\left|\textrm{GW}\left(t\right)\right\rangle+\left\langle\textrm{GW}\left(t\right)\right|\hat{H}\left|\textrm{GW}\left(t\right)\right\rangle\right]$. Variation with respect to the coefficients $C_{i,n}\left(t\right)$ yields the equation of motion:
$i\partial_{t}C_{i,n}\left(t\right)$ $=C_{i,n}\left(t\right)\left[\frac{U}{2}n\left(n-1\right)+M_{i}n\right]$
$-J\left[C_{i,n+1}\sqrt{n+1}\left(\sum_{j}\Phi_{j}^{*}\right)\right$. (5)
$\left.+C_{i,n-1}\sqrt{n}\left(\sum_{j}\Phi_{j}\right)\right]$
where $\sum_{j}\Phi_{j}=\sum_{j}\sum_{n=0}^{n_{\textrm{max}}-1}C_{j,n+1}C_{j,n}^{*}\sqrt{n+1}$ and $\sum_{j}$ denotes the sum over neighboring lattice points.
Time-dependent Gutzwiller method automatically conserves the expectation value total particle number $\left\langle\hat{N}\right\rangle$ as long as $\hat{H},\hat{N}]=0$. This can be illustrated by considering a generic Gutzwiller state at time $t\left|\textrm{GW}\left(t\right)\right\rangle$. After a time step $dt$, the state becomes:
$\left|\textrm{GW}\left(t+dt\right)\right\rangle=\left|\textrm{GW}\left(t\right)\right\rangle-idt\hat{H}\left|\textrm{GW}\left(t\right)\right\rangle$ (6)
The expectation value of the total particle number at the next time step is
$\left\langle\hat{N}\left(t+dt\right)\right\rangle$ $=\left\langle\textrm{GW}\left(t+dt\right)\right|\hat{N}\left|\textrm{GW}\left(t+dt\right)\right\rangle$
$=\left\langle\hat{N}\left(t\right)\right\rangle+idt\left\langle\textrm{GW}\left(t\right)\right|\left[\hat{H},\hat{N}\right]\left|\textrm{GW}\left(t\right)\right\rangle$ (7)
$+\mathcal{O}\left(dt^{2}\right).$
indicating $\left\langle\hat{N}\right\rangle$ is conserved during evolution as long as $\hat{H},\hat{N}]=0$.